Solve 2h+3f=5.25 and h+2f=3.00
Answer by unlockmath(1688)

Hello,
We can solve this a couple different ways. Let's do substitution. To do that let's get the h alone in the second equation.
2h+3f=5.25
h+2f=3.00
It will look like:
h=3.00-2f Subtract 2f from both sides. Now plug this into the first equation so it looks like this:
2(3.00-2f)+3f=5.25 Solve for f to get:
6-4f+3f=5.25 Combine like terms and Subtract 6 from both sides gets us:
-f=-.75 or
f=.75 Now plug .75 into either original equation to find h.
2h+3(.75)=5.25
2h+2.25=5.25 Subtract 2.25 from both sides to get:
2h=3 Divide both sides by 2 to get:
h=1.5 So now we have what h and f are. use these in the equations abouve and it will work out.
RJ Toftness
www.math-unlock.com